OpenStudy (idealist10):

What's the partial derivative of (x+2)(x-3)y with respect to x?

OpenStudy (idealist10):

@ganeshie8

OpenStudy (anonymous):

\[\begin{align*} \frac{\partial }{\partial x}\left[(x+2)(x-3)y\right]&=\frac{\partial }{\partial x}[x+2](x-3)y+(x+2)\frac{\partial }{\partial x}[x-3]y\\&~~~~~~~~+(x+2)(x-3)\frac{\partial }{\partial x}[y]\\\\ &=(x-3)y+(x+2)y+0\end{align*}\]

OpenStudy (idealist10):

So the answer is (2x-1)y, right?

OpenStudy (anonymous):

Yes

OpenStudy (idealist10):

Thanks.